OpenStudy (unofficialllyy):

What is the average velocity during this 11 s interval? Original question: What is the instantaneous velocity of a freely falling object 11 s after it is released from a position of rest? Answer I got: -110 m/s

OpenStudy (visarika23):

Avg velocity= total displacement / total time Total time =11 Total displacement = 1/2*10*11*11

OpenStudy (mathmale):

You now have a formula for velocity as a function of time. Use the average velocity formula (or slope formula) to determine the average velocity over the interval [0, 11]. The formula looks like this:\[\frac{ v(x _{2})-(t _{1} )}{ t _{2}-t _{1} }\]

OpenStudy (mathmale):

I always appreciate it when you share your work, so that I have something speicifc to which to respond.

OpenStudy (unofficialllyy):

I don't understand what numbers im supposed to plug in

OpenStudy (unofficialllyy):

@mathmale

OpenStudy (welshfella):

what valiue did you use for the acceleration due to gravity?

OpenStudy (unofficialllyy):

9.8 m/s^2

OpenStudy (welshfella):

OK instantaneous velocity after 11s = acceleration * time = 9.8 * 11 m/sec

OpenStudy (unofficialllyy):

107.8 m/s

OpenStudy (welshfella):

or did you use 10? because you have a value of 110?

OpenStudy (unofficialllyy):

Oh right 10, I forgot my assignment rounds it up

OpenStudy (welshfella):

yes 107.8 is more accurate . We are taking the downerds direction to be positive

OpenStudy (welshfella):

110 is you take 10 as acce;eration.

OpenStudy (unofficialllyy):

Right.

OpenStudy (unofficialllyy):

There's another part to this question if you could help me with it once we are finished with this part.

OpenStudy (welshfella):

To calculate the velocity use mathmale's formula

OpenStudy (welshfella):

* average velocity

OpenStudy (unofficialllyy):

Im not sure what to plug in for all of the parts though

OpenStudy (welshfella):

avg velocity = v2 - v2 110 - 0 ------ = -------- m / sec t2 - t1 11 - 0

OpenStudy (unofficialllyy):

Okay, so 10 m/s

OpenStudy (mathmale):

Let me ask you this; AT WHAT TIME DOES THE INTERVAL BEGIN? Why not represent this with zero (0)? Then, at what time does the interval end? Next, evalute the velocity at the left end point and again at the right endpoint. Again, the formula for average velocity here would be:\[\frac{ v(x _{2})-v(t _{1} )}{ t _{2}-t _{1} }\]

OpenStudy (mathmale):

When you see the words "average of a function," think: "I have to find two different values of that function, based upon the interval endpoints a and b. In this case your a=0 sec and your b=11 sec. Evaluate velocity v(t) at each of those times. Then use the formula\[\frac{ v(x _{2})-v(t _{1} )}{ t _{2}-t _{1} }\]

OpenStudy (mathmale):

to find the average velocity over the given interval.

OpenStudy (mathmale):

You have done this correctly. I hope my summary of the work involved in finding the average value of a function over a time interval helps you understand and remember this material.